Line l goes throught the point (5,–2) and is perpendicular to the line through (–2, 1) and (–4, –2). Find the slope of l. Match your result to the correct answer below.
A) - 2/3
B) - 3/2
C) 7/3
D) 0

Answer by Alan3354(21550)

Line l goes throught the point (5,–2) and is perpendicular to the line through (–2, 1) and (–4, –2). Find the slope of l. Match your result to the correct answer below.
A) - 2/3
B) - 3/2
C) 7/3
D) 0
-----------
Find the slope, m, of line through (–2,1) and (–4,–2).
m = diffy/diffx = (1+2)/(-2+4)
m = 3/2
---------
The slope of line 1 is the negative inverse of 3/2, = -2/3.
A) -2/3